# The Yōkai Learning Environment: Tracking Beliefs Over Space and Time

### v1

@@ 4 The Yōkai Learning Environment
The score in Equation 1 is only given when the game ends successfully, and sometimes can be negative due to hint misplacements. Agents, however, may learn that it is better to avoid grouping cards altogether to avoid penalties. To address this, we modify the reward function. If the game ends without a win, we give a small negative reward to discourage this strategy:
$R=\begin{cases}S&\text{if game won}\\ -d_{a}-(|C|-c_{a})-i_{h}&\text{otherwise.}\end{cases}$ (3)
where $c_{a}$ is the number of correctly formed clusters, $i_{h}$ is the number of incorrectly placed hints, and $d_{a}\in\{0,1\}$ indicates whether the game was ended early. Since rewards are sparse, we shaped only the training reward to accelerate learning. We additionally reward observing a previously unobserved card (+1+1), grouping a single colour, and placing a correct hint (both +1+1) or an incorrect hint (1-1).
5 Baselines, Experiments, and Results
To evaluate the spatio-temporal belief tracking capabilities of agents in the YLE, we explore five complementary research questions (RQs). We begin by comparing agents with different neural architectures to establish baselines, then progressively test their capacity for belief modelling, generalisation, and reasoning in increasingly complex scenarios:
RQ1
How well do different policies reason over their memory (zero-order ToM)?
RQ2
Does learning belief representations and explicit action encoding improve agents’ performance?
RQ3
Do agents learn belief‑tracking strategies that let them coordinate with new partners?
RQ4
Do agents represent beliefs in their internal activations?
RQ5
Can agents scale to higher-order ToM scenarios involving four players?
Setup
We use the two-player 3x3 YLE version, which already presents a significant challenge for RL agents. Agents are trained using MAPPO Yu et al. (2022) for $10^{9}$ steps and evaluated on 1,000 games, across three random seeds. The maximum episode length was $32$ for this setting. We report four metrics: return (R), success rate (SR), number of colours clustered (NC) and the rate of successfully ending early (SEE). SEE captures how often agents end the game early and still win – a decision that requires strong belief inference under uncertainty that comes with significant risk of failure. We tuned key hyperparameters for our models (e.g., encoder size, depth, etc.); Training details are in Appendix C.
5.1 RQ1a: Perfect memory is not sufficient to solve the YLE
Agent Memory R $\uparrow$ SR in % $\uparrow$ NC $\uparrow$ SEE in % $\uparrow$
Human - $5.7\pm 0.2$.5) $95.0\pm 2.5$.8) $3.0\pm 0.0$.0) $65.0\pm 4.8$.8)
GCN-M perfect $0.9\pm 0.0$.5) $26.8\pm 1.3$.4) $1.6\pm 0.0$.3) $0.0\pm 0.0$.0)
GATv2-M perfect $1.9\pm 0.0$.4) $58.2\pm 1.3$.5) $2.3\pm 0.0$.2) $0.0\pm 0.0$.0)
RM-M perfect $1.6\pm 0.1$.1) $50.7\pm 3.5$.0) $2.2\pm 0.0$.6) $0.0\pm 0.0$.0)
CNN-M perfect $2.3\pm 0.0$.3) $71.8\pm 0.8$.8) $2.6\pm 0.0$.1) $0.0\pm 0.0$.0)
CNN-LSTM implicit $0.2\pm 0.0$.1) $3.9\pm 0.4$.1) $0.8\pm 0.0$.3) $1.0\pm 0.1$.8)
CNN-GRU implicit $0.3\pm 0.0$.2) $2.3\pm 0.0$.3) $0.7\pm 0.0$.1) $2.1\pm 0.1$.8)
CNN-S5 implicit $0.1\pm 0.0$.1) $3.0\pm 0.3$.8) $0.7\pm 0.0$.2) $0.1\pm 0.0$.2)
CNN-TrXL explicit $1.5\pm 0.0$.3) $43.5\pm 0.9$.7) $2.1\pm 0.0$.2) $0.0\pm 0.0$.0)
ToMTrack explicit $1.6\pm 0.0$.3) $48.3\pm 1.0$.7) $2.2\pm 0.0$.2) $0.0\pm 0.0$.0)
Table 1: Agent performance on the 3x3 YLE. The first block compares encoder architectures with access to perfect memory (-M). The second block evaluates different temporal architectures with the best-performing CNN encoder. There, we compare implicit and explicit memory modules. Best results are in bold. ToMTrack (final block) is the best method overall when not given perfect memory.
Figure 3: Training curves for the 3x3 YLE, with bands showing the standard error over three seeds. Left: comparison of encoders with perfect memory. Right: comparison of memory architectures.
To evaluate if agents can reason about their own memory (zero-order ToM), we first compared different encoder architectures in a perfect memory setting: In this setting, agents were given perfect memory by including all previously observed cards in their observations. This experiment also served to set a performance upper bound. Given the YLEs spatial reasoning task and graph-based nature, we compared CNN-, GCN- Kipf and Welling (2016), GATv2- Brody et al. (2021), and Relation Module (RM)-based encoders Zambaldi et al. (2018). Architectural details are provided in Appendix C.2. To anchor results, we additionally establish a human performance baseline based on results from 25 games played by 5 participants (ages 24 - 28; see Appendix F).
As shown in Table 1 (top), all agents severely underperform the human baseline. Among agents, despite YLEs graph-based nature, the CNN encoder outperformed both GNNs and the RM in all metrics except SEE (see Figure 3, left). We hypothesise this is due to YLE’s dynamic spatial layout: card positions and relationships change frequently, which may be better captured by CNNs than by message-passing. In contrast to humans (SEE 65%), agents failed to solve the task reliably or finish early. This suggests that the YLE requires more than memory: It demands the ability to reason about asymmetric knowledge to maintain common ground over time and to act under uncertainty.
5.2 RQ1b: Explicit memory modules show stronger zero-order ToM than implicit ones
While the perfect memory setting establishes an upper bound, it assumes agents retain all past observations without limitation. To better reflect realistic constraints and to asses whether agents can reason over their own internal knowledge representations, we evaluated models with different temporal architectures. This tests their capacity to track and use information over time – a key aspect of zero-order ToM. We tested LSTM- Hochreiter and Schmidhuber (1997), GRU Cho et al. (2014), S5- Smith et al. (2023), and TransformerXL-based (Dai et al., 2019, TrXL) architectures, paired with the best-performing CNN encoder. S5 has recently excelled on memory tasks in Samsami et al. (2024), while TrXL includes an explicit memory mechanism shown effective in memory-intensive RL Parisotto et al. (2020).
As shown in Table 1 (middle) and Figure 3 (right), TrXL outperformed all other temporal models ($p<0.05$, one-sided paired t-test). Nonetheless, there is still a performance gap between TrXL and the best perfect memory baseline CNN-M. Contrary to expectations from prior work Samsami et al. (2024), S5 offered no clear advantage over traditional RNNs. LSTM, GRU, and S5 agents occasionally ended games early. This behaviour appears to reflect a form of gambling, as agents terminate the game after very few actions and without sufficient information-gathering. This suggests a failure to track beliefs over time, as agents that end early do so without sufficient inference – contrasting with human players, whose early-ending behaviour reflects confident ToM reasoning. This suggests a failure to reliably recall past observations, leaving guessing as the primary strategy. This tendency is further reflected in the short average game length ($22.9\pm 15.5$, $14.0\pm 15.6$ and $30.4\pm 1.42$, respectively). TrXL agents never finish early.
Taken together, these results indicate that explicit memory architectures, such as TrXL, are better suited for zero-order ToM reasoning. Nonetheless, the gap in performance relative to the perfect memory baseline demonstrates the difficulty of learning robust belief-tracking strategies in YLE.
5.3 RQ2: Modelling beliefs over cards and distilling partner policies improves results
Figure 4: Our ToMTrack agent generates beliefs over card colours using an auxiliary prediction head.
Figure 5: Left: compares return. Right: shows success rate. ToMTrack trains at least $1.5\times$ faster.
Given the partial observability of the YLE, agents must maintain a belief $b$ over the hidden colours $c$ of all cards. To support this, we designed ToMTrack, a method for learning an implicit model over cards’ colour (see Figure 4). ToMTrack is an online representation learning RL approach inspired by previous work on offline belief representation learning Wang et al. (2023). At each timestep $s_{t}$, the policy $\pi$ should learn useful representations that contain information about the cards’ colours, no matter if they have been previously observed or inferred (first-order ToM, recall Figure 1). We thus want our policy to model $p(c_{i}|s_{t}),\,\forall\,0<i<|Y|$. To achieve this, we added an auxiliary belief classification head $f_{\text{belief}}$ to the policy network. This head maps the intermediate representation learned by the policy to a distribution over possible card colours. During training, we used the ground truth card colours as labels. We used an auxiliary belief prediction task Jaderberg et al. (2017) to train $f_{\text{belief}}$. Specifically, we added a cross-entropy loss $\mathcal{L}^{\text{CE}}$ to the actor, $\mathcal{L^{\prime}}_{\text{a}}=\mathcal{L}_{\text{a}}+\lambda\mathcal{L}_{a}^{\text{CE}}$, and critic loss, $\mathcal{L^{\prime}}_{\text{c}}=\mathcal{L}_{\text{c}}+\lambda\mathcal{L}^{\text{CE}}_{c}$; where $\lambda\geq 0$ is a hyperparameter.
In addition, since information is often implicitly communicated using actions, we add a second input encoder to explicitly process the last action taken by the active agent. This enables the policy to explicitly interpret observed behaviour and update its beliefs about card colours accordingly, see Figure 1.
We also evaluated other baselines based on the simplified action decoder (SAD) Hu and Foerster (2020) or intrinsic motivation, but found that they underperformed, see Appendix H. In contrast, ToMTrack outperforms CNN-TrXL ($48.3$ vs $43.5$ SR) and decreases the gap to CNN-M ($71.8$ SR) (Table 1). These results demonstrate that modelling latent beliefs and interpreting others’ actions improves agent coordination.
5.4 RQ3: Agents do not exhibit transferable models of partner belief inference
Figure 6: Zero-shot coordination performance measured by SR. We report additional metrics in Appendix E.2. Self-play agents do not learn to coordinate with others. Other-Play agents (OP) have some success at partner generalisation. However, the additional complexity of YLE requires more symmetries as input to OP (+C+R) compared to Hanabi, which only requires +C.
Figure 7: ToM reasoning influences Refer to caption ’s decision. Without ToM reasoning, no move is optimal. A zero-order agent (memory-based) may move either the red or blue card, as both would group two cards. A first-order agent that correctly interprets Refer to caption ’s action (see Figure 1) will recognise that moving the blue card completes a full cluster. Agents are rewarded during training for forming complete clusters. An example of second-order ToM reasoning can be found in Appendix B.3.
Effective strategies for tracking one’s own and others’ beliefs should generalise to novel partners. We test this via zero-shot coordination (Hu et al., 2020, ZSC), where agents must cooperate with independently trained partners at test time. Success in ZSC indicates that agents do not rely on partner-specific conventions but instead learn to track beliefs, interpret actions, and reason about others’ knowledge in a generalisable way.
We first evaluate ZSC performance of the previously trained CNN-M and ToMTrack agents using cross-play (XP), where agents are paired with separately trained partners. Figure 6 shows that XP SR drops by $88.5\%$ on average, indicating a strong reliance on learned conventions.
To address this, we retrain agents using other-play (Hu et al., 2020, OP), which prevents overfitting to arbitrary conventions by exploiting environment symmetries. In Hanabi Bard et al. (2020), colour symmetry alone (OP+C) suffices for strong ZSC Hu et al. (2020). We find that this is not the case in YLE. As shown in Figure 6, while both CNN-M and ToMTrack trained with OP+C improve XP performance, a significant SP-XP performance gap remains. Unlike Hanabi, YLE introduces spatial-temporal complexity and thus additional symmetry classes (see Appendix C.4), e.g. agents could form conventions around locations. We therefore also test OP+C+R, using rotational symmetries to break spatial conventions. This improves XP further for CNN-M and ToMTrack. For CNN-M it nearly closes the SP-XP gap ($33.2$ vs $27.4$).
However, while OP+C+R improves XP, this comes at a significant gap in overall SR (e.g. $86.5$ vs $41.5$ SP in CNN-M), suggesting that non-OP agents rely heavily on spatial and colour conventions. When these are unavailable, coordination suffers – implying that belief tracking and action interpretation remain difficult to learn directly. This suggests that CNN-M and ToMTrack do not acquire general belief strategies out-of-the-box, though OP-trained variants may begin to.
Agent Memory T0 $\downarrow$ T1 $\downarrow$ Wrong $\uparrow$
CNN-M-OP+C+R Perfect $13.8\pm 0.0$.5) $86.5\pm 1.4$.2) $30.3\pm 0.3$.1)
ToMTrack-OP+C+R Explicit $12.6\pm 0.0$.9) $98.8\pm 2.8$.3) $28.2\pm 0.2$.7)
Table 2: Results from the diagnostic ToM test in Figure 7. We report the average rank (based on action probabilities) of zero-order (T0), first-order (T1), and incorrect actions (e.g., mismatched card colours). We find that agents fail to make zero- or first-order inferences in this simple test.
[…]

### v2

@@ 4 The Yōkai Learning Environment
The game returns the final score only at termination; the score can be negative due to incorrect hint placements. To discourage trivial strategies (e.g., never attempting to win to avoid hint penalties), we modify the terminal reward from Equation 1:
$R=\begin{cases}S&\text{if the game is won},\\ -d_{a}-(|C|-c_{a})-i_{h}&\text{otherwise},\end{cases}$ (4)
where $|C|$ is the number of colours, $c_{a}$ is the number of correct clusters, $i_{h}$ the number of incorrect hints (colour not included in the hint), and $d_{a}\in\{0,1\}$ indicates early termination. We additionally apply reward shaping to accelerate learning (rewarding newly observed cards and correct hints, both +1+1 each), which is annealed to zero during training and disabled at evaluation time.
Successful Early Ending (SEE)
Table 1: We evaluate steps-per-second (SPS) of YLE on a Nvidia H100 by taking random actions in $n$ parallel environments. YLE scales to hundreds of thousands of SPS.
Num Envs $n$ 512 1,024 2,048
2-player 9C 139,107 274,669 559,251
2-player 16C 135,662 144,636 148,269
Because unused hint cards yield the largest reward, every additional hint implicitly reduces the maximum attainable score, creating strong time pressure to end the game once sufficient common ground has been established. We therefore introduce Successful Early Ending (SEE) as an additional metric next to reward. SEE is the rate at which agents terminate the game early and still win. It is complementary to average game length: long games indicate overly cautious play, while short but unsuccessful games reflect guessing without sufficient belief formation. This positions SEE as a diagnostic signal of functional ToM reasoning.
Implementation
Our Jax-based implementation scales efficiently on GPU, reaching hundreds of thousands of steps per second (see Table 1) and performing comparably to other JAX-based RL environments (Rutherford et al., 2023; Nikulin et al., 2023; Matthews et al., 2024; Ruhdorfer et al., 2025b).
5 Adapting ZSC Methods to YLE
5.1 Other-Play in YLE
Other-Play (Hu et al., 2020) modifies self-play by optimizing expected return under random symmetry transformations drawn from the known symmetries of the Dec-POMDP. In practice, OP is implemented as domain randomization: at the start of each episode, each agent samples a transformation and interacts with a permuted view of the same underlying environment. This discourages convergence to arbitrary symmetry-breaking conventions that fail under cross-play.
In Hanabi, the dominant symmetry corresponds to permutations of card colours. Since card positions carry no semantic meaning and there is no spatial structure, enforcing invariance to colour permutations is sufficient to prevent most seed-dependent conventions.
In the YLE, symmetry structure is richer due to the spatial layout and card movement. Symmetry breaking can arise from both card colours and the spatial configuration. Enumerating all Dec-POMDP symmetries in complex environments is generally intractable in practice, which is a known limitation of OP. We therefore focus on the two most salient symmetry families that directly induce symmetry-breaking: (i) colour permutations and (ii) spatial grid rotations. Both admit symmetry breaking, e.g., clustering specific colours in fixed regions. We thus define the symmetry group as $\Phi=\Phi_{\text{col}}\times\Phi_{\text{rot}}$, where $\Phi_{\text{col}}$ permutes colours and $\Phi_{\text{rot}}=\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$ rotates the grid.
Unlike Hanabi, multiple symmetry classes exist in the YLE. Using only a subset may therefore leave residual symmetry-breaking conventions, which we show empirically in section 6.
5.2 Off-Belief Learning in YLE
Off-Belief Learning (Hu et al., 2021) constructs a sequence of policies and belief models. At level $\ell$, policy $\pi_{\ell}$ chooses actions that are optimal assuming (i) past actions were generated by a reference policy $\pi_{\ell-1}$, used to form beliefs, and (ii) future actions will follow $\pi_{\ell}$. Iterating this operator yields a hierarchy $\pi_{0}\rightarrow\pi_{1}\rightarrow\dots\rightarrow\pi_{L}$ that controls reasoning depth and, produces a unique policy per level given $\pi_{0}$. OBL relies on a learned belief model $\mathcal{B}_{\pi}$ that predicts the unknown state-action history $\tau_{t}$ given a joint policy $\pi$ and an agent’s action-observation history $\tau_{t}^{i}$, i.e. $\mathcal{B}_{\pi}(\tau_{t}|\tau_{t}^{i})$. In Hanabi, this belief predicts hidden card identities in the observing player’s hand.
In YLE, the (private, unknown) state consists of the unobserved colours of face-down Yokai cards. Accordingly, we train $\mathcal{B}_{\pi}$ to predict a posterior over the global board configuration conditioned on the agent’s AOH. In doing so, we closely follow the methodology of Hu et al. (2021): Our belief model features an LSTM (Hochreiter & Schmidhuber, 1997) encoder and autoregressive decoder that predicts a colour per card slot (i.e. nine slots for nine cards). Cards have a unique id associated with them in the observation which makes it possible for the belief model to map cards in observations to these slots. Per slot, we ensure that cards are sampled consistently with the ego agents AOH: We respect the global card colour restrictions (i.e. 9 cards of 3 colours) and only sample cards which are unknown first-hand to the agent from $\mathcal{B}$.
To our knowledge this is the first implementation of OBL outside of Hanabi. Implementing OBL outside of Hanabi highlights some limitations of OBL as formulated in Hu et al. (2021). First, hints in the YLE are not guaranteed truthful. The belief model must therefore treat hints as unreliable evidence, making posterior estimation more challenging than in Hanabi. Second, OBL theory requires a fictitious rollout length $K$ such that the acting agent moves again. While two-player Hanabi requires $K=2$, YLE’s multi-step turn structure requires $K=4|\mathcal{N}|$, increasing fictitious rollout cost substantially. In practice, this makes OBL expensive relative to other ZSC methods both in terms of additional environment steps and memory requirements. Such limitations are highlighted by testing ZSC algorithms in additional environments, a contribution of this work.
6 Experiments
We evaluate the YLE as a benchmark for zero-shot coordination by comparing self-play (SP) and cross-play (XP) performance across leading ZSC methods. We first show that the YLE is solvable in self-play, with agents achieving super-human performance in the simplest settings. However, strong self-play does not translate into strong cross-play. Across methods that achieve near-perfect inter-seed XP in Hanabi, we observe persistent SP–XP gaps and substantial degradation in XP performance in the YLE, indicating that existing ZSC approaches do not generalise to its demands.
6.1 Setup
Agents are trained using independent PPO (IPPO) (de Witt et al., 2020) for $10^{9}$ steps and evaluated on 5,000 games. Our default policy consists of a 4-layer CNN encoder (64 filters, kernel size $(3,3)$, ReLU, stride 1), followed by a linear projection, a GRU, and actor–critic heads. All hidden dimensions are 256. For OBL, we adopt the standard public–private architecture (Hu et al., 2021; Cui et al., 2022) with separate CNN encoders for private and public observations.
We optimise with Adam (Kingma & Ba, 2015) ($\epsilon{=}10^{-5}$), $\gamma{=}0.99$, $\lambda{=}0.95$, value loss coefficient $0.5$, four PPO epochs with four mini batches of 128 steps, and linearly annealed reward shaping. We tune the PPO entropy coefficient $\alpha$ per setting, sweeping $\alpha\in[0.01,0.08]$. We report results over 12 random seeds per configuration.
We report return (R), successful early ending (SEE), and game length (LEN). SEE measures how often agents terminate early and still win, and is decomposed into early-ending rate (EE) and calibration $\textbf{WEE}=P(\text{W}|\text{EE})$. Our primary comparison metric is cross-play return.
Maximum scores and episode lengths vary with the number of players and environment size (9C vs 16C) due to the differing number of available hints22244, 55, 66 for 2, 3, and 4 players respectively in 9C; $7$, 99, $10$ in 16C.. In 9C YLE, the maximum episode length is $32$, $40$, and $48$, and the theoretical maximum score is $20$, $25$, and $30$ for 2, 3, and 4 players respectively. Achieving this maximum requires immediate early termination without interacting with the board. By contrast, a policy that always uses all hints correctly and never terminates early obtains only $4$, 55, and $6$ respectively. Further details are provided in subsection C.2 in the supplementary.
6.2 Zero-Shot Coordination Experiments
Table 2: CNN-GRU SP and XP performance under different algorithms and settings in the YLE with humans (H). To isolate coordination challenges from memory limitations we use a perfect memory setting (M; indicated by ) where we show agents the values of already seen cards even if they have been moved. Full YLE is indicated by . Each entry in table is the mean $\pm$ standard error of the mean computed from $12$ independent seeds. The table shows $180$ trained policies. Best results are in bold. The column headers are: V = Card version (9C, 16C) and P = # of players (2, 3, 4). The difference between highest average XP and SP return is 7.54.8=2.77.5-4.8=2.7 (a $43.9\%$ difference).
Self-Play (SP) Cross-Play (XP)
Alg. V P M R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$ R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$
H 9C 2 $5.7\pm 0.0$.5) $65\pm 11.$ $65\pm 11.$ $100\pm 0.$
SP 9C 2 $7.5\pm 0.0$.3) $89\pm 7.$ $89\pm 8.$ $92\pm 8.$ $25.0\pm 0.0$.6) $1.6\pm 0.0$.2) $18\pm 3.$ $35\pm 5.$ $50\pm 3.$ $29.7\pm 0.0$.4)
HE 9C 2 $7.5\pm 0.0$.1) $95\pm 0.$ $6\pm 0. 99$$9\pm 0$. $24.6\pm 0.0$.1) $4.1\pm 0.0$.8) $50\pm 1.$ $68\pm 10.$ $69\pm 7.$ $27.2\pm 0.0$.7)
OP 9C 2 $6.6\pm 0.0$.2) $84\pm 2.$ $90\pm 2.$ $93\pm 1.$ $25.2\pm 0.0$.2) $4.8\pm 0.0$.1) $63\pm 2.$ $75\pm 2.$ $83\pm 1.$ $26.7\pm 0.0$.3)
OBL1 9C 2 $4.0\pm 0.0$.2) $68\pm 7.$ $76\pm 8.$ $83\pm 7.$ $28.0\pm 0.0$.4) $2.3\pm 0.0$.3) $43\pm 7.$ $56\pm 9.$ $76\pm 2.$ $29.4\pm 0.0$.4)
OBL2 9C 2 $4.3\pm 0.0$.2) $69\pm 7.$ $79\pm 7.$ $81\pm 7.$ $27.7\pm 0.0$.4) $2.1\pm 0.0$.2) $38\pm 4.$ $50\pm 5.$ $75\pm 2.$ $29.5\pm 0.0$.2)
OBL3 9C 2 $4.4\pm 0.0$.4) $73\pm 7.$ $82\pm 7.$ $81\pm 8.$ $27.3\pm 0.0$.4) $2.2\pm 0.0$.3) $41\pm 6.$ $55\pm 6.$ $73\pm 3.$ $29.1\pm 0.0$.4)
OBL4 9C 2 $4.5\pm 0.0$.2) $77\pm 3.$ $87\pm 3.$ $89\pm 2.$ $27.2\pm 0.0$.2) $2.8\pm 0.0$.3) $53\pm 8.$ $66\pm 8.$ $78\pm 3.$ $28.5\pm 0.0$.5)
OBL5 9C 2 $4.8\pm 0.0$.2) $80\pm 3.$ $89\pm 2.$ $90\pm 2.$ $27.0\pm 0.0$.2) $2.5\pm 0.0$.3) $47\pm 7.$ $61\pm 7.$ $75\pm 3.$ $28.9\pm 0.0$.3)
SP 9C 3 $7.8\pm 0.0$.6) $64\pm 10.$ $67\pm 11.$ $73\pm 12.$ $33.4\pm 0.1$.1) $0.2\pm 0.0$.1) $0\pm 0.$ $6\pm 1.$ $17\pm 1.$ $39.5\pm 0.0$.1)
HE 9C 3 $0.9\pm 0.0$.0) $3\pm 0.$ $100\pm 0.$ $3\pm 0.$ $1.0\pm 0.0$.0) $0.9\pm 0.0$.0) $4\pm 0.$ $100\pm 0.$ $4\pm 0.$ $1.0\pm 0.0$.0)
OP 9C 3 $3.0\pm 0.0$.8) $27\pm 0.$ $38\pm 11.$ $37\pm 11.$ $36.2\pm 0.1$.1) $1.2\pm 0.0$.2) $10\pm 1.$ $29\pm 4.$ $34\pm 3.$ $38.0\pm 0.0$.2)
SP 9C 4 $0.5\pm 0.0$.2) $3\pm 1.$ $50\pm 15.$ $3\pm 1.$ $38.1\pm 0.3$.0) $0.1\pm 0.0$.0) $1\pm 0.$ $64\pm 20.$ $1\pm 0.$ $36.6\pm 0.3$.7)
HE 9C 4 $0.7\pm 0.0$.1) $3\pm 0.$ $100\pm 0.$ $3\pm 0.$ $17.0\pm 0.3$.9) $0.7\pm 0.0$.0) $2\pm 0.$ $100\pm 0.$ $2\pm 0.$ $6.6\pm 0.2$.1)
OP 9C 4 $0.2\pm 0.0$.0) $1\pm 0.$ $49\pm 14.$ $1\pm 0.$ $38.8\pm 0.2$.6) $0.2\pm 0.0$ $1\pm 0.$ $68\pm 6.$ $2\pm 0.$ $36.5\pm 0.1$.5)
SP 16C 2 $0.0\pm 0.0$.0) $0\pm 0.$ $25\pm 12.$ $0\pm 0.$ $52.6\pm 0.1$.8) $0.0\pm 0.0$.0) $0\pm 0.$ $25\pm 16.$ $0\pm 0.$ $53.2\pm 0.1$.7)
As introduced, we compare SP, high entropy IPPO (HE) , OP, and OBL in a perfect-memory setting to isolate coordination challenges. For HE, we follow Forkel et al.: in 2P we use $\lambda_{\text{GAE}}=0.85$ and $\alpha=0.07$, and in multi-player settings we sweep $\alpha\in\{0.01,0.02,\dots,0.07\}$. For other 2P settings we use $\alpha=0.05$, while for 3P and 4P we use $\alpha=0.01$. We verify these choices via ablations below. Our final results are the result of training and comparing $400$ policies. To anchor results, we additionally establish a human performance baseline based on results from $25$ games played by $5$ participants (ages 24 - 28; see Appendix F). Our experiments run on NVIDIA V100, L40S, and H100 GPUs. A single seed takes between 6 and 12 hours depending on player and card count. Due to computational cost of training multiple levels and the methods performance, we evaluate OBL only in the 2P setting. We show XP matrices in Appendix H (Figures 15 - 24) and training curves in Appendix I which show our policies and belief models converging.
Table 2 highlights four key results. First, state-of-the-art ZSC methods exhibit persistent SP–XP gaps, indicating failure to learn fully symmetric policies in the YLE. Despite the large discrete action space in YLE, agents consistently achieve strong self-play performance, indicating that the main difficulty arises from coordination. Second, XP performance remains below the human baseline despite agents having perfect memory. Third, HE — the state-of-the-art method for Hanabi – underperforms OP in the YLE and collapses in multi-player settings, often learning to terminate immediately ($\textbf{LEN}=1.0$ and $\textbf{LEN}=3.8$). Increasing entropy makes coordination impossible before reaching high XP scores, suggesting that entropy tuning alone is insufficient for YLE. OBL also underperforms OP. Notably, this reverses the ranking observed in Hanabi, where HE outperforms OP. Fourth, difficulty increases systematically with additional players and in the 16C setting, all of which remain unsolved even w.r.t. SP returns. These results establish the YLE as a challenging benchmark for ZSC and cooperative MARL.
6.3 Ablations and Additional Experiments
Table 3: We sweep OP PPO entropy coefficients $\alpha$ to verify our experimental setting.
Self-Play (SP) Cross-Play (XP)
$\alpha$ V P M R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$ R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$
0.01 9C 2 $3.3\pm 0.0$.3) $32\pm 8.$ $41\pm 10.$ $46\pm 11.$ $29.3\pm 0.0$.6) $1.1\pm 0.0$.2) $12\pm 2.$ $28\pm 4.$ $47\pm 4.$ $30.2\pm 0.0$.4)
0.03 9C 2 $6.2\pm 0.0$.1) $85\pm 1.$ $92\pm 1.$ $92\pm 1.$ $25.6\pm 0.0$.2) $4.2\pm 0.0$.1) $56\pm 2.$ $76\pm 3.$ $74\pm 1.$ $26.7\pm 0.0$.3)
0.05 9C 2 $6.6\pm 0.0$.2) $84\pm 2.$ $90\pm 2.$ $93\pm 1.$ $25.2\pm 0.0$.2) $4.8\pm 0.0$.1) $63\pm 2.$ $75\pm 2.$ $83\pm 1.$ $26.7\pm 0.0$.3)
0.08 9C 2 $0.1\pm 0.0$.0) $\pm 0. 0±0$$\pm 0$. $0\pm 0.$ $32.0\pm 0.0$.0) $0.0\pm 0.0$.0) $\pm 0. 0±0$$\pm 0$. $0\pm 0.$ $32.0\pm 0.0$.0)
0.01 9C 3 $3.0\pm 0.0$.8) $27\pm 0.$ $38\pm 11.$ $37\pm 11.$ $36.2\pm 0.1$.1) $1.2\pm 0.0$.2) $10\pm 1.$ $29\pm 4.$ $34\pm 3.$ $38.0\pm 0.0$.2)
0.03 9C 3 $1.0\pm 0.0$.1) $8\pm 4.$ $62\pm 13.$ $11\pm 6.$ $19.5\pm 0.5$.3) $0.1\pm 0.0$.1) $1\pm 0.$ $44\pm 14.$ $1\pm 0.$ $24.9\pm 0.5$.0)
0.01 9C 4 $0.2\pm 0.0$.0) $1\pm 0.$ $49\pm 14.$ $1\pm 0.$ $38.8\pm 0.2$.6) $\pm 0.0 1±0.1\pm 0$. $68\pm 6.$ $2\pm 0.$ $36.5\pm 0.1$.5)
0.03 9C 4 $0.1\pm 0.0$.1) $1\pm 0.$ $17\pm 11.$ $1\pm 0.$ $42.5\pm 0.3$.9) $0.1\pm 0.0$.1) $1\pm 0.$ $33\pm 10.$ $1\pm 0.$ $38.7\pm 0.3$.0)
PPO Entropy Coefficient $\alpha$
The PPO entropy coefficient $\alpha$ plays an important role in ZSC. Both because it improves SP performance due to the added exploration empirically in the YLE but also since high entropy coefficients improve XP performance. Specifically, Forkel et al. (2026) show that in Hanabi, near-perfect inter-seed XP can be achieved by picking a high enough $\alpha$ that maximises XP performance while not yet degrading the policy due to the added randomness. In Table 3, we sweep $\alpha$ for OP. In the 2P setting, moderate entropy improves XP performance, consistent with prior findings. However, higher entropy does not achieve state-of-the-art YLE inter-seed XP performance and is outperformed by OP. In 3P and 4P settings, increasing entropy rapidly destabilises learning, leading to policy collapse before coordination benefits can materialise.
These results have two implications. First, even when combining symmetry-based training and entropy tuning, persistent SP–XP gaps remain in YLE. Second, unlike Hanabi, inter-seed XP in the YLE is not solvable via entropy threshold tuning alone.
Other-Play Symmetries
Table 4: OP with varying symmetry groups (G): full $\Phi$, only $\Phi_{\text{col}}$, and only $\Phi_{\text{rot}}$.
Self-Play (SP) Cross-Play (XP)
G V P M R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$ R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$
$\Phi$ 9C 2 $6.6\pm 0.0$.2) $84\pm 2.$ $90\pm 2.$ $93\pm 1.$ $25.2\pm 0.0$.2) $4.8\pm 0.0$.1) $63\pm 2.$ $75\pm 2.$ $83\pm 1.$ $26.7\pm 0.0$.3)
$\Phi_{\text{col}}$ 9C 2 $7.6\pm 0.0$.1) $97\pm 0.$ $97\pm 1.$ $100\pm 0.$ $24.7\pm 0.0$.1) $2.1\pm 0.0$.6) $24\pm 8.$ $39\pm 9.$ $55\pm 5.$ $29.5\pm 0.0$.7)
$\Phi_{\text{rot}}$ 9C 2 $6.4\pm 0.0$.3) $77\pm 7.$ $82\pm 7.$ $86\pm 8.$ $27.0\pm 0.0$.5) $4.8\pm 0.0$.2) $63\pm 3.$ $74\pm 3.$ $85\pm 2.$ $27.0\pm 0.0$.3)
As described in subsection 5.1, the YLE symmetry group factorises as $\Phi=\Phi_{\text{col}}\times\Phi_{\text{rot}}$. In Table 4, we compare OP trained with the full group $\Phi$ to ablations using only $\Phi_{\text{col}}$ or $\Phi_{\text{rot}}$. We find that using $\Phi$ and $\Phi_{\text{rot}}$ yields nearly identical XP performance, with $\Phi$ only slightly reducing game length. In contrast, $\Phi_{\text{col}}$ achieves strong SP but substantially lower XP, underperforming both $\Phi$ and $\Phi_{\text{rot}}$. This raises an interesting question for implementing OP in practice: How robust is OP to only partial knowledge of symmetry classes? We show that any single symmetry class alone may not necessarily sufficient for achieving good XP performance in the YLE.
Alternative Neural Architectures
Table 5: We compare several neural encoders (GATv2, GCN, and RM) and adding additional hidden layers (+1 HL, +2 HL). We find that our architecture is best among them.
Self-Play (SP)
Alg. V P M R $\uparrow$ SEE $\uparrow$ $E$ $\uparrow$ WEE $\uparrow$ LEN $\downarrow$
SP 9C 2 $7.5\pm 0.0$.3) $89\pm 7.$ $89\pm 8.$ $92\pm 8.$ $25.0\pm 0.0$.6)
GATv2 9C 2 $6.9\pm 0.0$.1) $87\pm 1.$ $87\pm 1.$ $100\pm 0.$ $26.0\pm 0.0$.3)
[…]